the Beta train is going 40km/h slower than the Lightning train, the lightning train makes the 850km journey in 3 hours less time. what is the speed of both trains?
Answer by richwmiller(17219)

r*t=d
B=L-40,
850=(t-3)*L,
850=t*B
B~~88.3205, L~~128.321, t~~9.62404